# Tracing the Techno-Supremacy Doctrine: A Critical Discourse Analysis of the AI Executive Elite

### v1

@@ 3.1 Data Analysis
This study adopts and adapts Norman Fairclough’s approach to Critical Discourse Analysis (CDA) (Fairclough, 2015). His framework rests on the fundamental assumption that discourse, understood as language as a form of social practice, is not only influenced by broader social structures (e.g., the media, economic systems, the state) but also contributes to constituting, reproducing, and changing these structures. His approach consists of three stages of analysis. The description stage involves identifying the values embodied in the formal linguistic properties of the text (e.g., vocabulary, grammar, structure). The interpretation stage aims to understand how the text was produced and could be interpreted to create meaning. Finally, the explanation stage focuses on why discourse is the way it is and what its effects are on the broader social context, particularly on power relations, ideologies, and social struggles.
@@
CDA has been critiqued primarily around three areas: potential bias given the analyst’s political positioning, lack of methodological rigor, and lack of empirical validation (cf. (O’Halloran, 2022; Stubbs, 1997; Toolan, 1997; Widdowson, 2008)). The methodology developed for this study was specifically designed to address these concerns. The overall goal was to enhance analytical rigor and transparency while retaining the critical power of Fairclough’s model. To mitigate the risk of analyst bias, the research design combines high transparency (e.g., in the selection of authors and texts) with critical self-reflexivity. In response to critiques of methodological rigor, a two-phase ‘T’ analysis was developed that includes deep qualitative investigations guided by patterns and hypotheses generated from a comprehensive coding of the entire corpus. The coding schemes developed provide a clear and consistent framework for identifying discursive features, thereby increasing the transparency and potential reproducibility of the analysis and mitigating the risk of cherry-picking data to fit a preconceived argument. Moreover, novel quantitative metrics were developed, derived directly from the coded linguistic data. This quantitative approach addresses critiques of CDA’s empirical grounding by providing a concrete, data-driven basis for identifying patterns, making comparisons between texts, and substantiating claims about the overall leanings and balance of the discourse. Finally, this study adopts a focused scope for the explanation stage to address the critique concerning the empirical validation of the text-context link. The societal effects of discourse and cognitive reception by audiences are notoriously difficult to establish without the use of extensive empirical methods, which are beyond the scope of this project. Therefore, the analysis refrains from making speculative claims about such impacts. Instead, the focus is on providing an explanatory analysis of potential causes, grounding explanations in an examination of the socio-historical context.
3.1.1 Broad Analysis
The first phase, represented by the top bar of the ‘T’, functions as an initial screening. It comprises a broad analysis across the corpus via relatively shallow versions of the CDA stages of description and interpretation, followed by a quantitative analysis to identify overarching patterns. The goal of this phase was to gain a detailed understanding of each author’s adherence to TSD (cf. objective O1) and their responses to AI-related concerns (cf. objective O2). First, upon an initial reading of the texts, the point (i.e., the summary interpretation of the text intended by the author) and the intended target audience (e.g., peers, potential followers, adversaries) for each text were identified. This analysis is considered a relatively shallow interpretation because the situational and intertextual contexts of each text were not analyzed, as prescribed by Fairclough. Second, on a closer reading, each text was thoroughly annotated with objective-specific codes (cf. Tables 1 and 2). Despite its thoroughness, this analysis is considered a relatively shallow description because the focus was on sentences rather than their linguistic features (e.g., vocabulary, structure, use of the passive voice). Finally, a statistical analysis was conducted, for which two metrics were developed (defined below). These metrics were used to analyze each text in isolation and to compare and contrast texts. Based on this quantitative analysis, specific phenomena were identified for further investigation in the second phase.
Code Name Description
@@ 3.1.1 Broad Analysis
TG-PE Prioritization of Efficiency An overemphasis on efficiency and convenience as the primary values driving the adoption and deployment of technology.
@@
TG-SO Superior Outcomes A belief that effectively addressing the technologically framed version of a problem achieves the optimal or superior outcome for the underlying complex societal issue.
@@
SL-SF Solution Focus A preoccupation with narrow, often technological solutions for complex problems.
@@
TC-AA Always the Answer An excessive conviction that technology, regardless of the context, is the optimal or primary solution to any problem.
@@
CT-DE Devaluation of Non-Tech A tendency to undervalue or disregard non-technological solutions, human wisdom, and social context in favor of technological interventions.
@@
CT-UF Utopian Future A belief in the transformative power of future technologies to create a better, even utopian world.
Note. Anti versions of these codes were used to annotate explicit contradictions of core TSD tenets (e.g., ANTI-CT-UF for CT-UF).
Table 1: TSD Presence and Reproduction Coding Scheme
Code Name Description
ACK-CR External Critique Acknowledging a concern from others.
@@
ADD-JU Justification Addressing a concern via justification or contextualization, arguing that, while possibly real, it is acceptable due to overriding benefits, necessity for progress, historical precedent, or context.
ADD-RE Refutation Addressing a concern via refutation or counter-argument against its validity or significance.
ADD-SN Non-Tech Solution Addressing a concern via a proposed solution or mitigation that is not primarily technology-based.
ADD-ST Tech Solution Addressing a concern via a proposed solution or mitigation primarily based on technology.
MAR-DE Deflection Avoiding addressing a concern directly by changing the subject, shifting blame, or attacking the critic.
@@
Table 2: Response to AI-Related Concerns Coding Scheme
Two metrics were developed as heuristics to guide the quantitative analysis: TSD Adherence (TDSA) and TSD Balance (TSDB). TSDA aims to synthesize the prevalence of pro- and anti-TSD discourse, as well as how responses to AI-related concerns might reinforce or counteract TSD within each analyzed text. Conversely, TSDB measures the balance between pro- and anti-TSD discourse within each text. Used together, these metrics enable a rich classification, allowing for a distinction between one-sided advocacy (high positive TSDA, low TSDB), one-sided critique (high negative TSDA, low TSDB), and texts characterized by significant internal debate (TSDA near zero, high TSDB). TSDA and TSDB are defined in terms of a pro- and an anti-TSD component, denoted by PRO-TSD and ANTI-TSD, respectively. TSDA is defined as their difference, while TSDB is defined as the percentage of the smaller with respect to the sum of both: [^1] If both components are the same but different than 0, TSDB is 0.5 (perfect balance). If only one is 0, TSDB is 0 (perfect one-sidedness). If both are 0, then TSDB is undefined, but this can only happen for irrelevant texts with no annotations.
Both PRO-TSD and ANTI-TSD are defined as a weighted sum of the normalized frequencies of specific codes. [^2] Frequencies are normalized to ensure text comparability by dividing raw counts by the text’s number of words and multiplying by 1,000 (i.e., code frequency per 1,000 words). Table 3 shows which codes contribute to which component to what extent. Note that codes are grouped into sub-components: TSD Core Expression (TCE), TSD-Reinforcing Responses (TRR), and their respective anti versions. The assignment of weights, while subjective, was a deliberate methodological choice grounded in the CDA principle of ideological centrality and discursive force. Specifically, greater weight was assigned to codes representing either core ideological tenets that form TSD’s foundation or forceful discursive strategies that directly confront or dismiss opposing views. Conversely, a lesser weight was given to statements that function as more specific applications, manifestations, or indirect strategies. This weighted approach is preferred over the alternative of equal weighting, which would make the equally subjective but theoretically weaker assumption that all discursive moves are ideologically equivalent.
[…]
PRO-TSD TCE CT-DE (Devaluation of Non-Tech) 2
@@
TRR ADD-ST (Tech Solution) 2
@@
ANTI-TSD ANTI-TCE $\alpha$ 2
$\beta$ 1
@@
$\alpha$. Anti codes for TCE codes with weight 2 (i.e., ANTI-CT-DE, ANTI-CT-MP, …).
$\beta$. Anti codes for TCE codes with weight 1 (i.e., ANTI-SL-SF, ANTI-SL-UI, …).
@@
3.1.2 Deep Analysis
@@
The second phase, represented by the vertical stem of the ‘T,’ comprises deeper qualitative investigations of selected phenomena via detailed description, interpretation, and/or explanation to understand their function and manifestation in detail. This phase aimed to inspect concrete phenomena identified in the previous phase more deeply through a qualitative lens, thereby complementing the previous, mainly quantitative analysis. It involved applying the full-fledged CDA stages of description, interpretation, and/or explanation.
The description stage involved analyzing linguistic features (e.g., vocabulary, grammar) for the values they embody: experiential, expressive, and/or relational. Experiential values are related to how the author represents the world. They are mainly tied to Tech Goggles and Solutionism (i.e., how the author frames problems and envisions potential solutions). Expressive values are related to the author’s evaluations and stances. They are mainly tied to Techno-Chauvinism and the Cult of Technology (i.e., what problem-solving approaches the author values and prefers). Finally, relational values are related to the relationships the author assumes, establishes, or negotiates with the intended audience.
The interpretation stage involved analyzing how the linguistic features identified during the description stage function communicatively to establish the text’s point and purpose. This stage required analyzing each text’s context in terms of production and interpretation conditions (situational context), and the historical series to which it belongs (intertextual context).
The explanation stage involved analyzing discursive strategies over time in the context of broader power struggles. The goal was to suggest potential explanations in terms of social determinants for why the discourse has changed in the way it has. According to Fairclough, this stage also involves analyzing the discourse in terms of its social effects. However, due to time, space, and methodological constraints, this study focuses solely on social determinants.
3.2 Data Collection
3.2.1 The AI Executive Elite
@@
This group comprises individuals with significant decision-making authority over the way AI is funded, developed, and deployed globally. First, the elite includes the CEOs and CTOs of the top five companies with respect to market share for generative AI models. As of writing, these are Microsoft (including Microsoft AI) (39%), Amazon (including AWS) (19%), Google (including Google Deepmind) (15%), OpenAI (9%), and Anthropic (4%) (Fernandez, 2025). Second, it includes Managing Partners (MPs) of venture capital firms in the top ten with respect to assets under management, which have also invested in the companies mentioned above that are exclusively devoted to AI (i.e., OpenAI and Anthropic). As of writing, these are Andreessen Horowitz, General Catalyst Partners, Sequoia Capital, and Khosla Ventures (noa, 2025a, b; Peck, 2025). Third, the elite includes the CEO and CTO of the leading provider of Graphics Processing Units, NVIDIA (Fernandez, 2025). Finally, the elite includes two individuals, Mark Zuckerberg and Elon Musk, due to their influential public profiles and substantial involvement in the AI field. While other notable individuals (e.g., chief scientists, prominent academic researchers, influential journalists) unquestionably contribute to shaping the AI discourse, by focusing specifically on C-suite executives and Managing Partners, this study intentionally prioritizes the analysis of ideological discourse linked to direct developmental and financial power. The cultural and scientific influence exercised by other actors is critical, but it represents a different form of power that is beyond the scope of this particular investigation.
The identification of the AI executive elite was conducted in March 2025 via Google Search, with all data rigorously verified against the corresponding official company and firm websites. The final group comprises 19 individuals: 11 CEOs, three CTOs, and five Managing Partners (cf. Table 4).
@@
3.2.2 Corpus
The corpus comprises publicly available texts in which elite members discuss the impact of technology or AI on society, including opportunities and concerns (cf. Table 5). Eligible texts include book chapters, op-eds, and blog posts (personal or corporate) of at least 750 words published since 2017. These criteria are intended to include texts in which members of the AI executive elite substantively express their views to allow for in-depth CDA, which is why shorter texts (e.g., tweets) were omitted. The year 2017 was chosen because it marks a pivotal year for generative AI, with the seminal paper Attention Is All You Need (Vaswani et al., 2017) laying the theoretical foundations for modern AI. Unless published as blog posts or op-eds, spoken texts were omitted as they tend to be less concise, more prone to vagueness, significantly longer, and more difficult to analyze due to non-verbal communication, all of which pose challenges given the time constraints of this research.
Texts were gathered in March 2025 using Google Search. The initial pool contained 32 texts. Surprisingly, some influential and vocal individuals (e.g., Hassabis, Musk) have not published any eligible texts, suggesting that their ideological influence is projected through different media (e.g., social media posts, interviews, podcasts, product demos). Given the need for in-depth analysis and time constraints, the final corpus was obtained by selecting two texts from authors with the most substantial engagement on the topic. Author engagement was measured in terms of the total number of words written ($w$) and the number of distinct publication dates ($d$): $E=(w^{\prime}+d^{\prime})/2$, where $w^{\prime}$ and $d^{\prime}$ are the min-max normalization of $\log(w)$ and $d$, respectively. [^3] This normalization is used to rescale values to a range between 0 and 1. The logarithmic transformation is applied to the number of words because the distribution is very skewed (i.e., some individuals have written only a few hundred words, while others have written several thousand). Only authors with an engagement score higher than 0.5 were considered for this study. While all eligible authors wrote at least two texts in the initial pool, most wrote more. The final two texts per author were selected by prioritizing type diversity (two different types (e.g., book chapter and op-ed) over a single one), topic specificity (AI over technology in general), and length (longer texts over shorter ones) in that order.
@@ 3.2.2 Corpus
Who will control the future of AI? (Altman, 2024a) altman-2024-control
The Intelligence Age (Altman, 2024b) altman-2024-intelligence
@@
Why AI Will Save the World (Andreessen, 2023a) andreessen-2023-save
The Techno-Optimist Manifesto (Andreessen, 2023b) andreessen-2023-manifesto
@@
3.3 Limitations
The study’s findings should be considered in light of several methodological limitations. First, the analysis unavoidably involved a degree of subjective interpretation. This limitation was mitigated with the development of a systematic and transparent methodology. Moreover, each text was consciously approached from positions of “engagement and estrangement” (Janks, 1997), aiming to balance empathetic understanding with critical scrutiny. Second, the quantitative metrics (TSDA and TSDB), developed as heuristics to identify broad discursive patterns, should not be considered absolute or objective measures of an individual’s beliefs. Their function is merely to guide and ground deeper qualitative analyses by providing comparative data across the corpus. Finally, the construction of the corpus is limited to relatively long written texts, excluding shorter formats (e.g., tweets) and spoken discourse (e.g., interviews, podcasts, keynotes), which may employ different rhetorical strategies. Moreover, the research relies exclusively on public-facing texts, which are often strategic pronouncements. Therefore, the analysis focuses on constructed ideological positions rather than private beliefs or motivations.
4 The Spectrum of TSD Discourse

### v2

@@ 3 The ‘T’ Approach to Critical Discourse Analysis
This paper adopts and adapts Norman Fairclough’s approach to Critical Discourse Analysis (CDA) (Fairclough, 2015). His framework rests on the fundamental assumption that discourse, understood as language as a form of social practice, is not only influenced by broader social structures (e.g., the media, economic systems, the state) but also contributes to constituting, reproducing, and changing these structures. His approach consists of analyzing (written or spoken) text in three stages. The description stage involves identifying the values embodied in the text’s formal linguistic properties (e.g., vocabulary, grammar). The interpretation stage aims to understand how the text was produced and could be interpreted to create meaning. Finally, the explanation stage focuses on why discourse is the way it is and what its effects might be on the broader social context, particularly on power relations, ideologies, and social struggles.
@@
CDA has been critiqued primarily around three areas: potential bias given the analyst’s political positioning; lack of methodological rigor; and lack of empirical validation (cf. (O’Halloran, 2022; Stubbs, 1997; Toolan, 1997; Widdowson, 2008)). The methodology developed for this paper is specifically designed to address these concerns. The overall goal was to enhance analytical rigor and transparency while retaining the critical power of Fairclough’s approach. To mitigate the risk of analyst bias, this research design combines high transparency (e.g., in the selection of authors and texts, in the annotation of texts) with critical self-reflexivity. Each text was consciously approached from positions of “engagement and estrangement” (Janks, 1997), aiming to balance empathetic understanding with critical scrutiny. In response to critiques of methodological rigor, a two-phase ‘T’ analysis was developed that includes deep qualitative investigations guided by patterns and hypotheses generated from a comprehensive coding of the entire corpus. The coding schemes developed provide a clear and consistent framework for identifying discursive features, thereby increasing the transparency and potential reproducibility of the analysis and mitigating the risk of cherry-picking data to fit a preconceived argument. Novel quantitative metrics were developed, derived directly from the coded linguistic data. This quantitative approach addresses critiques of CDA’s empirical grounding by providing a concrete, data-driven basis for identifying patterns, comparing texts, and substantiating claims about the overall leanings and balance of the discourse. Finally, this paper adopts a focused scope for the explanation stage to address the critique concerning the empirical validation of the text-context link. The societal effects of discourse and audience cognitive reception are notoriously difficult to establish without extensive empirical methods, which are beyond the scope of this project. Therefore, the analysis refrains from making speculative claims about such impacts. Instead, the focus is on providing an explanatory analysis of potential causes, grounding explanations in an examination of the socio-historical context.
3.1 Broad Analysis
TSD Component Code Name Description
@@ 3.1 Broad Analysis
Tech Goggles TG-PE Prioritization of Efficiency An overemphasis on efficiency and convenience as the primary values driving technology deployment and adoption.
@@
TG-SO Superior Outcomes A belief that addressing the technologically framed version of a problem effectively achieves the optimal or superior outcome for the underlying complex societal issue.
@@
Solutionism SL-SF Solution Focus A preference for narrow solutions to complex problems.
@@
Techno-Chauvinism TC-AA Always the Answer An excessive conviction that technology, regardless of the context, is the optimal or primary solution to any problem.
@@
Cult of Technology CT-DE Devaluation of Non-Tech A tendency to undervalue or disregard non-technological solutions, human wisdom, and social context in favor of technological interventions.
@@
CT-UF Utopian Future A belief in the transformative power of current or future technologies to create a better, even utopian world.
Note. Analogous ‘anti’ codes are used to annotate explicit contradictions of the corresponding aspect (e.g., ANTI-TG-PE).
Table 1: Coding Scheme for TSD Presence and Reproduction
Response Type Code Name Description
Acknowledging ACK-CR External Critique Acknowledging a concern from others.
@@
Addressing ADD-JU Justification Addressing a concern, arguing that while possibly real, it is acceptable due to overriding benefits, necessity for progress, historical precedent, or context.
ADD-RE Refutation Addressing a concern via a counter-argument against its validity.
ADD-SN Non-Tech Solution Addressing a concern via a proposed solution or mitigation that is not primarily based on technology.
ADD-ST Tech Solution Addressing a concern via a proposed solution or mitigation that is primarily based on technology.
Marginalizing MAR-DE Deflection Avoiding addressing a concern directly by changing the subject, shifting blame, or attacking the critic.
@@
Table 2: Coding Scheme for Response to AI-Related Concerns
(Sub-) Components Codes Weight
PRO-TSD TSD Core Expression (TCE) CT-DE (Devaluation of Non-Tech) 2
@@
TSD-Reinforcing Responses (TRR) ADD-ST (Tech Solution) 2
@@
ANTI-TSD ANTI-TCE Anti TCE codes with weight 2 2
Anti TCE codes with weight 1 1
@@
Note. The assignment of weights is a deliberate methodological choice grounded in the CDA principle of ideological centrality and discursive force. Specifically, greater weight is assigned to codes representing either core ideological tenets that form TSD’s foundation or forceful discursive strategies that directly confront or dismiss opposing views. Conversely, a lesser weight is given to statements that function as more specific applications, manifestations, or indirect strategies.
@@
The first phase, represented by the top bar of the ‘T,’ functions as an initial screening. It comprises a broad analysis across the corpus using relatively shallow versions of the CDA stages of description and interpretation, followed by a quantitative analysis to identify patterns for investigation in the second phase. The goal of this phase is to gain a detailed understanding of each author’s adherence to TSD (cf. objective I) and their responses to AI-related concerns (cf. objective II).
This phase is composed of three steps. First, upon an initial reading, each text’s point (i.e., the summary interpretation of the text intended by the author) and its likely intended target audience (e.g., peers, potential followers, adversaries) are identified. Second, on a closer reading, individual sentences are annotated with custom codes corresponding to objectives I (cf. Table 1) and II (cf. Table 2), respectively. For example, the sentence “Technology will usher in an era of prosperity for all mankind.” would be annotated with codes TC-PD (Progress Driver) and CT-UF (Utopian Future). Finally, a quantitative analysis of code counts is conducted to determine the extent to which each text is pro- or anti-TSD. Each text is analyzed both in isolation and with respect to other texts to identify concrete phenomena to be qualitatively investigated in the second phase of the analysis.
Two metrics were developed for the quantitative analysis: TSD Adherence (TDSA) and TSD Balance (TSDB). TSDA aims to synthesize the prevalence of pro- and anti-TSD discourse and how responses to AI-related concerns might reinforce or counteract TSD within each analyzed text. Conversely, TSDB measures the balance between pro- and anti-TSD discourse within each text. Used together, these metrics enable a rich classification, allowing for a distinction between one-sided advocacy (high positive TSDA, low TSDB), one-sided critique (high negative TSDA, low TSDB), and texts characterized by significant internal debate (TSDA near zero, high TSDB). These metrics were developed as heuristics to identify broad discursive patterns and should not be considered absolute or objective measures of an individual’s beliefs. Their function is merely to guide and ground deeper qualitative analyses by providing comparative data across the corpus.
TSDA and TSDB are defined in terms of pro- and anti-TSD components, denoted PRO-TSD and ANTI-TSD, respectively. TSDA is defined as their difference, while TSDB is defined as the percentage of the smaller with respect to the sum of both:
PRO-TSD and ANTI-TSD are defined as a weighted sum of the normalized counts of specific codes according to Table 3.111Counts are normalized to ensure text comparability by dividing raw counts by the text’s number of words and multiplying by 1,000. For example, suppose a particular 2,000-word text has 6 sentences annotated with TC-PD and 4 with TG-PE. By normalizing counts and multiplying by the corresponding weights, we have that $\text{TCE}=((6/2000*1000)*2)+((4/2000*1000)*1)=8$. Suppose that TRR = 2, ANTI-TCE = 1, and ANTI-TRR = 4. Then, $\text{PRO-TSD}=\text{TCE}+\text{TRR}=10$ and $\text{ANTI-TSD}=\text{ANTI-TCE}+\text{ANTI-TRR}=5$. Therefore, $\text{TSDA}=5$ and $\text{TSDB}=0.\overline{3}$. According to these values, this text is pro-TSD ($\text{TSDA}>0$) and presents its stance in a moderately balanced way ($\text{TSDB}>0.25$). [^2] Perfect balance ($\text{TSDB}=0.5$) is achieved if both components are the same (but different than 0), while perfect one-sidedness ($\text{TSDB}=0$) is achieved if only one is 0. If both are 0, then TSDB is undefined, but this can only happen for irrelevant texts with no annotations.
3.2 Deep Analysis
@@
The second phase of the methodology, represented by the vertical stem of the ‘T,’ comprises deeper investigations of selected phenomena to understand their functions and manifestations in detail. This phase aims to inspect concrete phenomena identified in the previous phase more deeply and in context through a qualitative lens, thereby complementing the previous, mainly quantitative analysis. It involves applying the full-fledged CDA stages of description, interpretation, and/or explanation.
The description stage involves analyzing linguistic features for the values they embody: experiential, expressive, and/or relational. Experiential values are related to how the author represents the world (i.e., how the author frames problems and envisions potential solutions). These values are mainly tied to Tech Goggles and Solutionism. Expressive values are related to the author’s evaluations and stances (i.e., what problem-solving approaches the author values and prefers). These values are mainly tied to Techno-Chauvinism and the Cult of Technology. Finally, relational values correspond to the relationships the author assumes, establishes, or negotiates with the intended audience. The interpretation stage involves analyzing how the linguistic features identified during the description stage function communicatively to establish the text’s point and purpose. This stage requires analyzing each text’s situational and intertextual contexts (i.e., its production and interpretation conditions, and the historical series to which it belongs). Finally, the explanation stage involves analyzing discursive strategies over time in the context of broader power struggles. The goal is to suggest potential explanations in terms of social determinants for why the discourse has changed in the way it has (cf. objective III). [^3] This stage also involves analyzing the discourse in terms of its social effects; however, due to time, space, and methodological constraints, this paper focuses solely on social determinants. Though presented sequentially, these stages are intended to be iterative. Fairclough suggests that the analyst move between them to build a holistic, mutually explanatory account.
4 Empirical Study
@@
The empirical study consisted of applying the ‘T’ CDA methodology described previously to a corpus of publicly available texts authored by members of the AI executive elite. This elite comprises individuals with significant decision-making authority over the way AI is funded, developed, and deployed globally. It is defined by four criteria. First, the CEOs and CTOs of the top five companies by market share in generative AI models. As of writing, these are Microsoft (including Microsoft AI) (39%), Amazon (including AWS) (19%), Google (including Google Deepmind) (15%), OpenAI (9%), and Anthropic (4%) (Fernandez, 2025). Second, the Managing Partners (MPs) of the top ten venture capital firms by assets under management, provided they have invested in the aforementioned companies exclusively devoted to AI (i.e., OpenAI and Anthropic). As of writing, these are Andreessen Horowitz, General Catalyst Partners, Sequoia Capital, and Khosla Ventures (9; 46; L. Peck (2025)). Third, the CEO and CTO of the leading provider of Graphics Processing Units, Nvidia (Fernandez, 2025). Finally, two individuals, Mark Zuckerberg and Elon Musk, due to their influential public profiles and substantial involvement in the AI field. Elite members were identified in March 2025 via Google Search, with all data rigorously verified against the corresponding official company and firm websites. The final group comprises 19 individuals: 11 CEOs, three CTOs, and five MPs (cf. Table 4).
@@
Note. While other notable individuals (e.g., chief scientists, prominent academic researchers, influential journalists) unquestionably contribute to shaping the AI discourse, by focusing specifically on C-suite executives and MPs, this paper intentionally prioritizes the analysis of ideological discourse linked to direct developmental and financial power. The cultural and scientific influence exerted by other actors is critical, but it represents a different form of power beyond the scope of this investigation.
@@
Who will control the future of AI? (Altman, 2024b) altman-2024-control
The Intelligence Age (Altman, 2024a) altman-2024-intelligence
@@
Why AI Will Save the World (Andreessen, 2023b) andreessen-2023-save
The Techno-Optimist Manifesto (Andreessen, 2023a) andreessen-2023-manifesto
@@
The corpus contains texts in which elite members discuss the impact of technology or AI on society, including opportunities and concerns. It was constructed systematically in accordance with concrete criteria to avoid cherry-picking. Eligible texts are book chapters, op-eds, and blog posts (personal or corporate) of at least 750 words published since 2017. These criteria are intended to include texts in which members of the AI executive elite substantively express their views, allowing for in-depth CDA, which is why shorter texts (e.g., tweets) were omitted. The year 2017 was chosen because it marks a pivotal year for generative AI, with the seminal paper Attention Is All You Need (Vaswani et al., 2017) laying the theoretical foundations for modern AI. Unless published as blog posts or op-eds, spoken texts were omitted as they tend to be less concise, more prone to vagueness, significantly longer, and more difficult to analyze due to non-verbal communication, all of which pose challenges given the time constraints of this research.
Texts were gathered in March 2025 using Google Search. The initial pool contained 32 texts. Given the need for in-depth analysis and time constraints, the final corpus was selected from two texts by authors with the most substantial engagement on the topic. Author engagement was measured in terms of the total number of words written ($w$) and the number of distinct publication dates ($d$) as $E=(w^{\prime}+d^{\prime})/2$, where $w^{\prime}$ and $d^{\prime}$ are the min-max normalization of $\log(w)$ and $d$, respectively. [^4] This normalization is used to rescale values to a range between 0 and 1. The logarithmic transformation is applied to the number of words because the distribution is very skewed (i.e., some individuals have written only a few hundred words, while others have written several thousand). Only authors with an engagement score higher than 0.5 were considered for this research. While all eligible authors wrote at least two texts in the initial pool, most wrote more. The final two texts per author were selected by prioritizing text type diversity (e.g., a blog post and an op-ed are preferred over two blog posts), topic specificity (i.e., AI over technology in general), and length (i.e., longer texts over shorter ones) in that order.
The final corpus is shown in Table 5. Surprisingly, some influential and vocal individuals (e.g., Hassabis, Musk) have not published any eligible texts, suggesting that their ideological influence is projected through different media (e.g., social media posts, interviews, podcasts, product demos).
4.1 The Spectrum of TSD Discourse